# MuCo-KGC: Multi-Context-Aware Knowledge Graph Completion

### v1

@@ 1 Introduction
  • Description-Independence: Unlike prior approaches, MuCo-KGC does not require entity descriptions, making it applicable to knowledge graphs with sparse or missing metadata.
@@ 2 Related Work
Embedding-based techniques (KGE) capture a graph’s underlying semantics by representing entities and relationships as continuous vectors. These can further categorize into translation distance-based methods and semantic or textual matching-based methods. The first group includes TransE[2], RotatE [12], and others approaches [9], which train a scoring function to measure the distance between entities, modeling interactions as translation operations. In contrast, semantic or textual matching methods, such as DistMult [21], rely on scoring functions that compute the similarity between entities and relationships. These methods identify the most likely tail entity during testing by evaluating scores across all entities that share a specific head and relationship.
@@
Recent hybrid methods seek to combine the strengths of large language models (LLMs), prompt engineering [20] and traditional knowledge graph embedding (KGE) techniques. For example, Pretrain-KGC [25] integrates LLMs with classical models like TransE and RotatE by initializing embeddings in a shared vector space, subsequently refining them through iterative optimization to improve loss detection. Similarly, DIFT [7] employs fine-tuned LLaMA-2-7B models augmented with a knowledge adaptation module, enabling enhanced reasoning over KG-specific domains. Despite achieving state-of-the-art results, DIFT’s dependence on computationally intensive LLMs like LLaMA-2-7B limits its practicality in resource-constrained settings.
@@
A critical limitation of existing KGC methods is their insufficient exploitation of the structural properties of knowledge graphs. Many approaches overlook the rich contextual information embedded in the graph’s structure, such as neighboring entities and relationships tied to a given head entity and query relation. Key structural components that can provide valuable context include head context-entities (neighboring entities), head context-relations (neighboring relations), and the relationship context (information specific to the query relation).

### v2

@@ Abstract
00footnotetext: Accepted in PAKDD DSFA Special Session
@@ 1 Introduction
  • Description-Independence: Unlike prior approaches, MuCo-KGC does not require entity descriptions, making it applicable to KG’s with sparse or missing metadata.
@@ 2 Related Work
KGE capture a graph’s underlying semantics by representing entities and relationships as continuous vectors. These can further categorize into translation distance-based methods and semantic or textual matching-based methods. The first group includes TransE[9], RotatE [10], and others approaches [11], which train a scoring function to measure the distance between entities, modeling interactions as translation operations. In contrast, semantic or textual matching methods, such as DistMult [5], rely on scoring functions that compute the similarity between entities and relationships. These methods identify the most likely tail entity during testing by evaluating scores across all entities that share a specific head and relationship.
@@
Recent hybrid methods seek to combine the strengths of large language models (LLMs), prompt engineering [13] and traditional KGE techniques. For example, Pretrain-KGC [7] integrates LLMs with classical models like TransE and RotatE by initializing embeddings in a shared vector space, subsequently refining them through iterative optimization to improve loss detection. Similarly, DIFT [8] employs fine-tuned LLaMA-2-7B models augmented with a knowledge adaptation module, enabling enhanced reasoning over KG-specific domains. Despite achieving state-of-the-art results, DIFT’s dependence on computationally intensive LLMs like LLaMA-2-7B limits its practicality in resource-constrained settings.
@@
A critical limitation of existing KGC methods is their insufficient exploitation of the structural properties of KG’s. Many approaches overlook the rich contextual information embedded in the graph’s structure, such as neighboring entities and relationships tied to a given head entity and query relation. Key structural components that can provide valuable context include head context-entities (neighboring entities), head context-relations (neighboring relations), and the relationship context (information specific to the query relation).